# Mixing, segregation, and collapse transitions of interacting copolymer rings

### v2

@@ 5.3 The SC-M phase boundary
Next, we consider the phase diagram in the region $\beta_{c}>\beta_{c}^{*}$ and $\beta_{m}>\beta_{m}^{*}$. We performed simulations crossing phase boundaries as shown in the right panel of figure 4 by keeping fixed the parameter $\beta_{c}$ (mutual attractions) at values $\{0.4,0.5,0.75\}$ inside the SC phase while varying $\beta_{m}$ across SC-M phase boundaries. A similar approach was followed to examine the M-SC boundary by varying $\beta_{c}$ while keeping $\beta_{m}$ fixed at the values $\{0.4,0.5,0.75\}$, following the corresponding arrows in the left panel of figure 4.
@@
Similar to the SE-M phase boundary (see section 5.2) the graphs in figure 9 shows the presence of finite-size phase behaviour at a critical value of $\beta_{m}$. The location of the phase boundary can be estimated by again examining crossings between the curves for the variance and $d_{cm}$ in figure 9. These are plotted in figure 10 (the left panel is the data determined from $Var(k_{m})$ and the right panel corresponds to the results from $d_{cm}$). The data for $\beta_{c}=0,0.2,0.3$ are also included for reference in the graphs (see figure 8) – these are the empty symbols. Notice that these empty symbols coincide within statistical uncertainty for different values of $\beta_{c}=0,0.2,0.3$, but that when $\beta_{c}>\beta_{c}^{*}$ then the data (the filled-in symbols) spread to larger values of $\beta_{m}$. This trend is seen in both the left and right panels of figure 9. In figure 10, for $\beta_{c}=0.50$, the data extrapolate (purple triangles for $\beta_{c}=0.50$) to $\beta_{m}^{*}\approx 0.5$ when $\beta_{c}=0.5$, while for $\beta_{c}=0.75$ the data (red circles) seem to extrapolate to $\beta_{m}^{*}\approx 0.75$. The case $\beta_{c}=0.4$ (yellow/green triangles) is more delicate since it is closer to the line of $\theta$-points along the SE-SC phase boundary separating the SE-M and SC-M phase boundaries. The results in the right panel apparently extrapolate to a value $\beta_{m}^{*}\approx 0.4$ when $\beta_{c}=0.4$.
@@ 5.4 The M-SC phase boundary
Estimates of the location of the M-SC phase boundary, based on the crossings of the $Var(k_{c})$ curves and the $d_{cm}$ curves are plotted in figure 12. These points are again plotted against $1/\sqrt{n}$, and the vertical bar on the $y$-axis is the confidence interval of estimate of $\beta_{m}^{*}$ in reference [12] when $\beta_{c}=0$.
@@ 6 Topological entanglement
Since their computation would be prohibitively costly when the number of crossings $n_{c}$ after a planar projection is very large, we first simplify the geometrical entanglement of the system while keeping its topology unaltered [13, 15]. This is achieved by shrinking stochastically the polygons with the BFACF algorithm, simulated at a very small step fugacity $K$ [4, 1, 2]. The BFACF moves have the property of preserving the topology and, if $K$ is sufficiently small, the system will often eventually shrink to its minimal length compatible with its topology. Even if the minimal length is not always reached, the shrunk configuration, once projected onto a plane, displays, most of the time, a sufficiently small number of crossings (well below $50$) to dramatically reduce the time required to compute the topological invariants [13, 22].
Figure 15: Topological entanglement as the SE-M phase boundary is crossed by varying $\beta_{m}$: Linking probability (top row) and knotting probability (bottom row) as a function of the mutual attraction parameter $\beta_{m}$ for three fixed values of the self attraction parameter $\beta_{c}$, (a) $0.00$, (b) $0.20$ and (c) $0.28$. Note that $\beta_{c}=0.28$ is very close to the estimate of the $\theta$-point $\beta_{c}^{*}$ (where the collapse transitionoccurs) when $\beta_{m}=0$.
@@ 6.1 Fixing $\beta_{m}$ and varying $\beta_{c}$
In figure 13 the linking and knotting properties of the model are explored for $\beta_{m}\in\{0,0.2,0.3\}$ so that the SE-SC phase boundary is crossed when $\beta_{c}$ increases beyond the critical point $\beta_{c}^{*}\approx 0.28$. In the SE phase (the top row of panels in the figure for $\beta_{c}<\beta_{c}^{*}$) the components are linked with positive probability, increasing as the length of the polygons increases and as $\beta_{m}$ approaches from below the transition value $\beta_{m}^{*}$. Increasing $\beta_{c}$ reduces the probability of linking to zero in the SC phase. This behaviour is expected, as the two components are segregated and compact in this phase, excluding each other and so making linking unlikely. In contrast, in the SE phase, the two polygons, although segregated, are also expanded, allowing strands of one to enter the other and to become entangled, giving rise to a positive probability of topological linking between the two components.
@@ 7 Discussion
In section 2, we describe the model in more detail and make comments about the phase diagram. We provide some rigorous results in section 3, where we prove existence of a phase boundary between the segregated phases and the mixed phase. The two parts of this boundary are the SE-M and SC-M boundaries in figure 2. In addition we argue that the phase boundary between the SE and SC phases is a line at constant $\beta_{c}$.

### v3

@@ 5.3 The SC-M phase boundary
Next, we consider the phase diagram in the region $\beta_{c}>\beta_{c}^{0}$ and $\beta_{m}>\beta_{m}^{*}(0)$. We performed simulations crossing phase boundaries as shown in the right panel of figure 4 by keeping fixed the parameter $\beta_{c}$ (mutual attractions) at values $\{0.4,0.5,0.75\}$ inside the SC phase while varying $\beta_{m}$ across SC-M phase boundaries. A similar approach was followed to examine the M-SC boundary by varying $\beta_{c}$ while keeping $\beta_{m}$ fixed at the values $\{0.4,0.5,0.75\}$, following the corresponding arrows in the left panel of figure 4.
@@
Similar to the SE-M phase boundary (see section 5.2) the graphs in figure 9 shows the presence of finite-size phase behaviour at a critical value of $\beta_{m}$. The location of the phase boundary can be estimated by again examining crossings between the curves for the variance and $d_{cm}$ in figure 9. These are plotted in figure 10 (the left panel is the data determined from $Var(k_{m})$ and the right panel corresponds to the results from $d_{cm}$). The data for $\beta_{c}=0,0.2,0.3$ are also included for reference in the graphs (see figure 8) – these are the empty symbols. Notice that these empty symbols coincide within statistical uncertainty for different values of $\beta_{c}=0,0.2,0.3$, but that when $\beta_{c}>\beta_{c}^{0}$ then the data (the filled-in symbols) spread to larger values of $\beta_{m}$. This trend is seen in both the left and right panels of figure 9. In figure 10, for $\beta_{c}=0.5$, the data extrapolate (purple triangles for $\beta_{c}=0.5$) to $\beta_{m}^{*}(0.5)\approx 0.5$, while for $\beta_{c}=0.75$ the data (red circles) seem to extrapolate to $\beta_{m}^{*}(0.75)\approx 0.75$. The case $\beta_{c}=0.4$ (yellow/green triangles) is more delicate since it is closer to the line of $\theta$-points along the SE-SC phase boundary separating the SE-M and SC-M phase boundaries. The results in the right panel apparently extrapolating to a value $\beta_{m}^{*}(0.4)\approx 0.4$ when $\beta_{c}=0.4$.
@@ 5.4 The M-SC phase boundary
Estimates of the location of the M-SC phase boundary, based on peaks in the $Var(k_{c})$ curves and the crossings between $d_{cm}$ curves are plotted in figure 12. These points are again plotted against $1/\sqrt{n}$, and the vertical bar on the $y$-axis is the confidence interval of estimate of $\beta_{m}^{*}$ in reference [12] when $\beta_{c}=0$.
@@ 6 Topological entanglement
Since their computation would be prohibitively costly when the number of crossings $n_{c}$ after a planar projection is very large, we first simplify the geometrical entanglement of the system while keeping its topology unaltered [13]. This is achieved by shrinking stochastically the polygons with the BFACF algorithm, simulated at a very small step fugacity $K$ [4, 1, 2]. The BFACF moves have the property of preserving the topology and, if $K$ is sufficiently small, the system will often eventually shrink to its minimal length compatible with its topology. Even if the minimal length is not always reached, the shrunk configuration, once projected onto a plane, displays, most of the time, a sufficiently small number of crossings (well below $50$) to dramatically reduce the time required to compute the topological invariants [13, 21].
Figure 15: Topological entanglement as the SE-M phase boundary is crossed by varying $\beta_{m}$: Linking probability (top row) and knotting probability (bottom row) as a function of the mutual attraction parameter $\beta_{m}$ for three fixed values of the self attraction parameter $\beta_{c}$, (a) $0.00$, (b) $0.20$ and (c) $0.28$. Note that $\beta_{c}=0.28$ is very close to the estimate of the $\theta$-point $\beta_{c}^{0}$ (where the collapse transitionoccurs) when $\beta_{m}=0$.
@@ 6.1 Fixing $\beta_{m}$ and varying $\beta_{c}$
In figure 13 the linking and knotting properties of the model are explored for $\beta_{m}\in\{0,0.2,0.3\}$ so that the SE-SC phase boundary is crossed when $\beta_{c}$ increases beyond the critical point $\beta_{c}^{0}\approx 0.28$. In the SE phase (the top row of panels in the figure for $\beta_{c}<\beta_{c}^{0}$) the components are linked with positive probability, increasing as the length of the polygons increases and as $\beta_{m}$ approaches from below the transition value $\beta_{m}^{*}(\beta_{c})$. Increasing $\beta_{c}$ reduces the probability of linking to zero in the SC phase. This behaviour is expected, as the two components are segregated and compact in this phase, excluding each other and so making linking unlikely. In contrast, in the SE phase, the two polygons, although segregated, are also expanded, allowing strands of one to enter the other and to become entangled, giving rise to a positive probability of topological linking between the two components.
@@ 7 Discussion
In section 2, we describe the model in more detail and make comments about the phase diagram. We provide some rigorous results in section 3, where we prove existence of a phase boundary between the segregated phases and the mixed phase. The two parts of this boundary are the SE-M and SC-M boundaries in figure 2. In addition we argue that the phase boundary between the SE and SC phases is a line at constant $\beta_{c}=\beta_{c}^{0}$.